OpenStudy (anonymous):

What is 28% of 75? A. 47 B. 28 C. 37 D. 21

OpenStudy (anonymous):

What do you think?

OpenStudy (mathstudent55):

In math, "of" usually means "times". Also, % means "divided by 100". That means 28% of 75 means 28/100 * 75 28/100 = 0.28, so you need to multiply 0.28 * 75

OpenStudy (anonymous):

75 x 28 / 100 = 21

OpenStudy (campbell_st):

just use \[75 \times 28 \div 100\] a percentage of a quatity is always \[quantity \times percentage \div 100\]

OpenStudy (briellshaw):

the answer is D divide 75 by 28%

OpenStudy (ashchu117):

here http://www.onlineconversion.com/percentcalc.htm

OpenStudy (ashchu117):

\.3./